Particles A and B carrying charge 3.0nC are at the base corners of an equilateral triangle 2.8 m on a side

What is the potential[relative to zero at infinity] at the apex of the triangle if particle B is replaced by a particle carrying a charge of -3.0nC?
1
Expert's answer
2021-10-13T16:56:22-0400

Since the triangle is equilateral, the electrostatic potential produced at the apex by the particle A is given as follows:


"\phi_A = k\dfrac{q_A}{r}"

where "k" is Coulomb constant, "q_A" is the charge of A, and "r" is the side of the triangle.

Similarly, the electrostatic potential produced at the apex by the particle B is given as follows:


"\phi_B = k\dfrac{q_B}{r}"

But, since in the later case "q_A = -q_B", the total potential will be:


"\phi = \phi_A + \phi_B = k\dfrac{q_A}{r} + k\dfrac{q_B}{r} = k\dfrac{q_A}{r} - k\dfrac{q_A}{r} = 0"

Answer. 0.